# MagicTailor: Component-Controllable Personalization in Text-to-Image Diffusion Models

### v1

@@ Abstract
Recent advancements in text-to-image (T2I) diffusion models have enabled the creation of high-quality images from text prompts, but they still struggle to generate images with precise control over specific visual concepts. Existing approaches can replicate a given concept by learning from reference images, yet they lack the flexibility for fine-grained customization of the individual component within the concept. In this paper, we introduce component-controllable personalization, a novel task that pushes the boundaries of T2I models by allowing users to reconfigure specific components when personalizing visual concepts. This task is particularly challenging due to two primary obstacles: semantic pollution, where unwanted visual elements corrupt the personalized concept, and semantic imbalance, which causes disproportionate learning of the concept and component. To overcome these challenges, we design MagicTailor, an innovative framework that leverages Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics and Dual-Stream Balancing (DS-Bal) to establish a balanced learning paradigm for desired visual semantics. Extensive comparisons, ablations, and analyses demonstrate that MagicTailor not only excels in this challenging task but also holds significant promise for practical applications, paving the way for more nuanced and creative image generation.
@@
Figure 1: (a) Illustration of personalization, demonstrating how text-to-image (T2I) diffusion models can learn and reproduce a visual concept from given reference images. (b) Illustration of component-controllable personalization, depicting a newly formulated task that aims to modify a specific component of a visual concept during personalization. (c) Example images generated by MagicTailor, showcasing the effectiveness of the proposed MagicTailor, a novel framework that adapts T2I diffusion models for component-controllable personalization. For clarity, the red and blue circles are used to highlight the target concept and component, respectively.
@@ 1 Introduction
Text-to-image (T2I) diffusion models (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; Chen et al., 2023) demonstrate remarkable capabilities in generating high-quality visual content from textual descriptions. These models are able to create images that closely match the provided prompts. But when certain visual concepts are difficult to articulate through natural language, they may face difficulties in accurately incorporating such elements into the generated images. To address this limitation, some approaches (Gal et al., 2022; Ruiz et al., 2023) enable T2I models to learn specific concepts from a few reference images, thus allowing for more faithful integration of these concepts into the generated images. This process, illustrated in Figure 1(a), is referred as personalization. However, existing personalization methods are limited to replicating predefined concepts and lack the capability to offer flexible, fine-grained control over these elements. This constraint significantly limits their practical applicability in real-world scenarios. A concept often comprises multiple components, such as a house consisting of walls, windows, and doors. Therefore, a more sophisticated challenge in personalization is determining how to effectively control and manipulate these individual components during the personalization process.
@@
This paper introduces a new task, component-controllable personalization, which aims to reconfigure specific components of personalized concepts using additional visual references (see Figure 1(b)). In this task, a T2I model is fine-tuned with reference images and corresponding category labels, allowing it to learn and generate the desired concept along with the given component. Achieving this capability would not only enable users to refine and customize concepts with precise control but also foster innovation and creativity, paving the way for novel ideas, inventions, and artworks across various creative domains.
@@
A straightforward approach to this task is to treat each component as a separate concept and use existing personalization methods to combine multiple concepts with suitable text prompts. However, this naive strategy falls short in component-controllable personalization due to the inherent complexity of handling visual semantics during learning. One of the key challenges in this task is semantic pollution (see Figure 2(a)), where undesired visual semantics inadvertently appear in generated images, thereby “polluting” the personalized concept. This occurs because the T2I model tends to blend visual semantics from different regions during the learning process. Simply masking out unwanted visual elements in reference images is not a viable solution, as it disrupts the overall visual context and leads to unintended compositions. Another significant challenge is semantic imbalance (see Figure 2(b)), which causes the T2I model to focus disproportionately on certain aspects, resulting in unfaithful personalization. This issue arises from the semantic disparity between the concept and component, highlighting the need for an effective learning paradigm to better manage concept-level (e.g., person) and component-level (e.g., hair) visual semantics.
@@
To address these challenges, we present MagicTailor, a novel framework that enables component-controllable personalization for T2I models (see Figure 1(c)). As shown in Figure 3, we first employ a text-guided image segmenter to generate segmentation masks for both the concept and component. Then, we introduce a technique called Dynamic Masked Degradation (DM-Deg), which transforms the original reference images into randomly degraded versions, dynamically perturbing undesired visual semantics. This approach helps suppress the model’s sensitivity to irrelevant visual details while preserving the overall visual context, thereby effectively mitigating semantic pollution. Next, we initiate a warm-up phase for the T2I model by jointly training it on these degraded images, using a masked diffusion loss to focus on the desired visual semantics and an attention loss to strengthen the correlation between these semantics and pseudo-words. To tackle the issue of semantic imbalance, we employ Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm designed for balancing the learning of visual semantics, to launch the second phase. In this paradigm, the online denoising U-Net performs sample-wise min-max optimization, while the momentum denoising U-Net applies selective preserving regularization. This balanced approach ensures more faithful and accurate personalization of the target concept and component.
@@
We validate the effectiveness of MagicTailor through comprehensive qualitative and quantitative experiments, demonstrating that it can achieve state-of-the-art (SOTA) performance in component-controllable personalization. Detailed ablation studies further confirm the impact of the key techniques integrated into MagicTailor. Additionally, we showcase its potential to enable a variety of further applications. In summary, the main contributions of this work are as follows:
  • We introduce a new task called component-controllable personalization for T2I models, enabling precise control over the individual components of concepts during personalization. Moreover, semantic pollution and semantic imbalance are identified as key challenges in this task.
  • We propose MagicTailor, a novel framework specifically designed for component-controllable personalization. This framework incorporates Dynamic Masked Degradation (DM-Deg) to dynamically perturb undesired visual semantics, and Dual-Stream Balancing (DS-Bal) to ensure balanced learning of visual semantics.
  • Comprehensive comparisons demonstrate that MagicTailor achieves superior performance in this task. Additionally, ablation studies and further applications highlight the effectiveness and versatility of the proposed method.
@@ 2 Related Works
Figure 2: Major challenges in component-controllable personalization. (a) Semantic pollution: (i) Undesired visual elements may inadvertently disturb the personalized concept. (ii) A simple mask-out strategy is ineffective and causes unintended compositions, whereas (iii) our DM-Deg effectively suppresses unwanted visual semantics, preventing such pollution. (b) Semantic imbalance: (i) Simultaneously learning the concept and component can lead to imbalance, resulting in concept or component distortion (here we present a case for the former). (ii) Our DS-Bal ensures balanced learning, enhancing personalization performance.
Text-to-Image Generation. Text-to-image (T2I) generation has made remarkable advancements in recent years, enabling the synthesis of vivid and diverse imagery based on textual descriptions. Early methods employed Generative Adversarial Networks (GANs) (Reed et al., 2016; Xu et al., 2018; Qiao et al., 2019; Zhu et al., 2019), and transformers (Ding et al., 2021; Ramesh et al., 2021; Ding et al., 2022; Yu et al., 2022; Bai et al., 2024) began to show the potential in conditioned image generation. More recently, the advent of diffusion models has ushered in a new era in T2I generation (Li et al., 2024; Saharia et al., 2022; Ramesh et al., 2022; Chen et al., 2023; Xue et al., 2024). Leveraging these models, a range of related applications has rapidly emerged, including image editing (Li et al., 2024; Mou et al., 2024; Huang et al., 2024; Feng et al., 2024; Bai et al., 2023), image completion and translation (Xie et al., 2023b; a; Lin et al., 2024), and controllable generation (Zhang et al., 2023; Wang et al., 2024b; Zheng et al., 2023). Despite advancements in T2I diffusion models, generating images that accurately capture specific, user-defined concepts remains challenging. This study explores component-controllable personalization, enabling precise adjustment of concept’s components through visual references.
Personalization. Personalization seeks to adapt T2I models to generate specific concepts using reference images. Initial approaches such as textual inversion (Gal et al., 2022) and DreamBooth (Ruiz et al., 2023) addressed this task by either optimizing a text embedding or fine-tuning the entire T2I model. Additionally, low-rank adaptation (LoRA) (Hu et al., 2021) has been widely adopted by the research community for personalization (Ryu, 2022), offering an efficient and lightweight solution. The scope of personalization has further expanded to accommodate multiple concepts (Kumari et al., 2023; Avrahami et al., 2023; Liu et al., 2023; Gu et al., 2024; Han et al., 2023; Gu et al., 2024). Besides, a growing body of works has explored tuning-free approaches to personalization (Xiao et al., 2023; Li et al., 2023; Shi et al., 2023; Wei et al., 2023; Wang et al., 2024a). However, these methods often rely on training an encoder with extensive domain-specific image datasets. There is also a category of works studying training-free schemes (Jeong et al., 2024; Zhang et al., 2024), but they generally suffer from inferior performance and tortuous inference processes. In light of that, Our MagicTailor goes with a widely-adopted paradigm of test-time optimization to achieve stable performance and precise control over the concept during personalization.
@@ 3 Methodology
Let ={($\mathcal{I}=\{(\{I_{nk}\}_{k=1}^{K},c_{n})\}_{n=1}^{N}$ represents a concept-component pair with $N$ samples of concepts and components, where each sample contains $K$ reference images $\{I_{nk}\}_{k=1}^{K}$ with a category label $c_{n}$. In this work, we focus on handling one concept and one component to make the task setting more practical. Specifically, we set $N=2$ and define the first sample as a concept (e.g., dog) while the second one as a component (e.g., ear). In addition, these samples are associated with the pseudo-words $\mathcal{P}=\{p_{n}\}_{n=1}^{N}$ serving as their text identifiers. The objective of component-controllable personalization is to fine-tune a text-to-image (T2I) model to accurately learn both the concept and component from $\mathcal{I}$. Using text prompts with $\mathcal{P}$, the fine-tuned model is expected to generate images that contain the personalized concept integrated with specified components.
In this section, we begin by providing an overview of the MagicTailor pipeline (refer to Section 3.1). Following this, we delve into its two core techniques: Dynamic Masked Degradation (DM-Deg, see Section 3.2) and Dual-Stream Balancing (DS-Bal, see Section 3.3).
Figure 3: Pipeline overview of MagicTailor. Using reference images as the inputs, MagicTailor fine-tunes a T2I diffusion model to learn both the target concept and component, enabling the generation of images that seamlessly integrate the component into the concept. Two key techniques, Dynamic Masked Degradation (DM-Deg, see Section 3.2) and Dual-Stream Balancing (DS-Bal, see Section 3.3), address the challenges of semantic pollution and semantic imbalance, respectively. For clarity, only one image per concept/component is presented and the warm-up stage is not depicted.
@@ 3.1 Overall Pipeline
The overall pipeline of MagicTailor is illustrated in Figure 3. The process begins with identifying the desired concept or component within each reference image $I_{nk}$, employing an off-the-shelf text-guided image segmenter to generate a segmentation mask $M_{nk}$ based on $I_{nk}$ and its associated category label $c_{n}$. Conditioned on $M_{nk}$, we introduce Dynamic Masked Degradation (DM-Deg) to perturb undesired visual semantics within $I_{nk}$, addressing semantic pollution. At each training step, DM-Deg transforms $I_{nk}$ into a randomly degraded image $\hat{I}_{nk}$, with the degradation intensity being dynamically regulated. Subsequently, these degraded images, along with structured text prompts, are used to fine-tune a T2I diffusion model to facilitate concept and component learning. The model is formally expressed as $\{\epsilon_{\theta},\tau_{\theta},\mathcal{E},\mathcal{D}\}$, where $\epsilon_{\theta}$ represents the denoising U-Net, $\tau_{\theta}$ is the text encoder, and $\mathcal{E}$ and $\mathcal{D}$ denote the image encoder and decoder, respectively. To promote the learning of the desired visual semantics, we employ the masked diffusion loss, which is defined as
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{n,k,\epsilon,t}\Big{[}\big{\|}\epsilon\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]},$ (1)
where $\epsilon\sim\mathcal{N}(0,1)$ is the unscaled noise, $z_{nk}^{(t)}$ is the noisy latent image of $\hat{I}_{nk}$ with a random time step $t$, $e_{n}$ is the text embedding of the corresponding text prompt, and $M_{nk}^{\prime}$ is downsampled from $M_{nk}$ to match the shape of $\epsilon$ and $z_{nk}$. Additionally, we also incorporate the cross-attention loss to strengthen the correlation between desired visual semantics and their corresponding pseudo-words, formulated as
@@
when $A_{\theta}(p_{n},z_{nk}^{(t)})$ is the cross-attention maps between the pseudo-word $p_{n}$ and the noisy latent image $z_{nk}^{(t)}$ and $M_{nk}^{\prime\prime}$ is downsampled from $M_{nk}$ to match the shape of $A_{\theta}(p_{n},z_{nk}^{(t)})$. Using $\mathcal{L}_{\text{diff}}$ and $\mathcal{L}_{\text{attn}}$, we first warm up the T2I model by jointly learning all samples, aiming to preliminarily inject the knowledge of visual semantics. The loss of the warm-up stage is defined as
@@
where $\lambda_{\text{attn}}=0.01$ is the loss weight for $\mathcal{L}_{\text{attn}}$. For efficient fine-tuning, we only train the denoising U-Net $\epsilon_{\theta}$ in a low-rank adaptation (LoRA) (Hu et al., 2021) manner and the text embedding of the pseudo-words $\mathcal{P}$, keeping the others frozen. Thereafter, we employ Dual-Stream Balancing (DS-Bal) to establish a dual-stream learning paradigm to address the challenge called semantic imbalance. In this paradigm, the online denoising U-Net $\epsilon_{\theta}$ conducts sample-wise min-max optimization for the hardest-to-learn sample, and meanwhile the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ applies selective preserving regularization for the other sample.

### v2

@@ Abstract
Recent text-to-image models generate high-quality images from text prompts but lack precise control over specific components within visual concepts. Therefore, we introduce component-controllable personalization, a new task that allows users to customize and reconfigure individual components within concepts. This task faces two challenges: semantic pollution, where undesirable elements distort the concept, and semantic imbalance, which leads to disproportionate learning of the target concept and component. To address these, we design MagicTailor, a framework that uses Dynamic Masked Degradation to adaptively perturb unwanted visual semantics and Dual-Stream Balancing for more balanced learning of desired visual semantics. The experimental results show that MagicTailor outperforms existing methods in this task and enables more personalized, nuanced, and creative image generation.
Figure 1: (a) Personalization Illustration: Showing how text-to-image diffusion models learn and generate a visual concept from reference images. (b) Component-Controllable Personalization Illustration: Introducing a task that aims to modify specific components of visual concepts during personalization. (c) Generated Images by MagicTailor: Demonstrating MagicTailor’s effectiveness in component-controllable personalization, with red and blue circles highlighting the target concept and component, respectively.
@@ 1 Introduction
Text-to-image (T2I) diffusion models [35, 39, 31, 5] have shown impressive capabilities in generating high-quality images from textual descriptions. While these models can generate images that align well with provided prompts, they struggle when certain visual concepts are hard to express in natural language. To address this, methods like [11, 36] enable T2I models to learn specific concepts from a few reference images, allowing for more accurate integration of those concepts into the generated images. This process, as shown in Fig. 1 (a), is referred as personalization.
However, existing personalization methods are limited to replicating predefined concepts and lack flexible and fine-grained control over their components. This limitation hinders their practical use in real-world applications. Since concepts often consist of multiple components, such as a house with walls, windows, and doors, a key problem in personalization is how to effectively control and manipulate these individual components.
Figure 2: Major challenges in component-controllable personalization. (a) Semantic pollution: (i) Undesired elements may interfere with the personalized concept. (ii) A simple mask-out strategy causes unintended results, while (iii) DM-Deg effectively suppresses unwanted semantics. (b) Semantic imbalance: (i) Simultaneously learning the concept and component can distort either one. (ii) DS-Bal ensures balanced learning, improving personalization. (c) Identity fidelity performance: Calculating DreamSim [10] scores on a collected dataset, we show that DM-Deg and DS-Bal can address these challenges for faithful generation.
@@
In this paper, we introduce component-controllable personalization, a new task that enables the reconfiguration of specific components within personalized concepts using additional visual references (Fig. 1 (b)). In this approach, a T2I model is fine-tuned with reference images and corresponding category labels, allowing it to learn and generate the desired concept along with the given component. This capability empowers users to refine and customize concepts with precise control, fostering creativity and innovation across various domains, from artworks to inventions.
@@
One key challenge of this task is semantic pollution (Fig. 2 (a)), where unwanted visual elements inadvertently appear in the generated image, “polluting” the personalized concept. This happens because the T2I model often mixes visual semantics from different regions during training. Masking out unwanted elements in reference images doesn’t solve the problem, as it disrupts the visual context and creates unintended compositions. Another challenge is semantic imbalance (Fig. 2 (b)), where the model overemphasizes certain aspects, leading to unfaithful personalization. This occurs due to the semantic disparity between the concept and component, necessitating a more balanced learning approach to manage concept-level (e.g., person) and component-level (e.g., hair) semantics.
Figure 3: Pipeline overview of MagicTailor. This method fine-tunes a T2I diffusion model using reference images to learn both the target concept and component, enabling the generation of images that seamlessly integrate the component into the concept. Two key techniques, Dynamic Masked Degradation (DM-Deg, see Sec. 3.2) and Dual-Stream Balancing (DS-Bal, see Sec. 3.3), address semantic pollution and semantic imbalance, respectively. For clarity, only one image per concept/component is shown, and the warm-up stage is omitted.
@@
To address these challenges, we present MagicTailor, a novel framework that enables component-controllable personalization for T2I models ( Fig. 1 (c)). We first use a text-guided image segmenter to generate segmentation masks for both the concept and component and then design Dynamic Masked Degradation (DM-Deg) to transform reference images into randomly degraded versions, perturbing undesired visual semantics. This method helps suppress the model’s sensitivity to irrelevant details while preserving the overall visual context, effectively mitigating semantic pollution. Next, we initiate a warm-up phase for the T2I model, training it on the degraded images using a masked diffusion loss to focus on the desired semantics and a cross-attention loss to strengthen the correlation between these semantics and pseudo-words. To address semantic imbalance, we develop Dual-Stream Balancing (DS-Bal), a dual-stream learning paradigm that balances the learning of visual semantics. In this phase, the online denoising U-Net performs sample-wise min-max optimization, while the momentum denoising U-Net applies selective preservation regularization. This ensures more accurate and faithful personalization of the target concept and component.
@@
We validate the effectiveness of MagicTailor through various qualitative and quantitative experiments, demonstrating state-of-the-art (SOTA) performance in component-controllable personalization. Detailed ablation studies further confirm the impact of the key techniques integrated into MagicTailor. Additionally, we show its potential for enabling a wide range of creative applications.
@@ 2 Related Works
Text-to-Image Generation.
Text-to-image (T2I) generation has made remarkable advancements in recent years, enabling the synthesis of vivid and diverse imagery based on textual descriptions. Early methods employed Generative Adversarial Networks (GANs) [33, 53, 28, 59], and transformers [6, 30, 7, 55, 4] also showed the potential in conditional generation. The advent of diffusion models has ushered in a new era in T2I generation [20, 39, 31, 5, 54]. Leveraging these models, a range of related applications has rapidly emerged, including image editing [20, 24, 16, 9, 3], image completion and translation [50, 49, 22], and controllable image generation [56, 45, 58]. Despite advancements in T2I diffusion models, generating images that accurately reflect specific, user-defined concepts remains a challenge. This study explores component-controllable personalization, which allows precise adjustment of specific concepts’ components using visual references.
Personalization.
Personalization seeks to adapt T2I models to generate given concepts using reference images. Initial approaches [11, 36] addressed this task by either optimizing text embeddings or fine-tuning the entire T2I model. Additionally, low-rank adaptation (LoRA) [15] has been widely adopted by the research community for personalization [37], offering an efficient and lightweight solution. The scope of personalization has expanded to encompass multiple concepts [19, 2, 13, 12, 25]. Besides, several studies [48, 21, 40, 46, 44] have explored tuning-free approaches for personalization, but these often necessitate training an encoder on large, domain-specific image datasets [57]. In this work, MagicTailor adopts the widely used paradigm of test-time optimization, providing stable performance and precise control over personalization.
@@ 3 Methodology
Let ={($\mathcal{I}=\{(\{I_{nk}\}_{k=1}^{K},c_{n})\}_{n=1}^{N}$ represents a concept-component pair with $N$ samples of concepts and components, where each sample contains $K$ reference images $\{I_{nk}\}_{k=1}^{K}$ with a category label $c_{n}$. In this work, we focus on a practical setting involving one concept and one component. Specifically, we set $N=2$ and define the first sample as a concept (e.g., dog) while the second one as a component (e.g., ear). In addition, these samples are associated with the pseudo-words $\mathcal{P}=\{p_{n}\}_{n=1}^{N}$ serving as their text identifiers. The goal of component-controllable personalization is to fine-tune a text-to-image (T2I) model to accurately learn both the concept and the component from $\mathcal{I}$. Using text prompts with $\mathcal{P}$, the fine-tuned model should generate images that integrate the personalized concept with the specified component.
This section begins by providing an overview of the MagicTailor pipeline in Sec. 3.1 and then delves into its two core techniques: Dynamic Masked Degradation in Sec. 3.2 and Dual-Stream Balancing in Sec. 3.3.
@@ 3.1 Overall Pipeline
The overall pipeline of MagicTailor is illustrated in Fig. 3. The process begins with identifying the desired concept or component within each reference image $I_{nk}$, employing an off-the-shelf text-guided image segmenter to generate a segmentation mask $M_{nk}$ based on $I_{nk}$ and its associated category label $c_{n}$. Conditioned on $M_{nk}$, we design Dynamic Masked Degradation (DM-Deg) to perturb undesired visual semantics within $I_{nk}$, addressing semantic pollution. At each training step, DM-Deg transforms $I_{nk}$ into a randomly degraded image $\hat{I}_{nk}$, with the degradation intensity being dynamically regulated. Subsequently, these degraded images, along with structured text prompts, are used to fine-tune a T2I diffusion model to facilitate concept and component learning. The model is formally expressed as $\{\epsilon_{\theta},\tau_{\theta},\mathcal{E},\mathcal{D}\}$, where $\epsilon_{\theta}$ represents the denoising U-Net, $\tau_{\theta}$ is the text encoder, and $\mathcal{E}$ and $\mathcal{D}$ denote the image encoder and decoder, respectively. To promote the learning of the desired visual semantics, we employ the masked diffusion loss, which is defined as:
$\mathcal{L}_{\text{diff}}\ =\ \mathbb{E}_{n,k,\epsilon,t}\Big{[}\big{\|}\epsilon_{n}\odot M_{nk}^{\prime}-\epsilon_{\theta}(z_{nk}^{(t)},t,e_{n})\odot M_{nk}^{\prime}\big{\|}_{2}^{2}\Big{]}\ ,$ (1)
where $\epsilon_{n}\sim\mathcal{N}(0,1)$ is the unscaled noise, $z_{nk}^{(t)}$ is the noisy latent image of $\hat{I}_{nk}$ with a random time step $t$, $e_{n}$ is the text embedding of the corresponding text prompt, and $M_{nk}^{\prime}$ is downsampled from $M_{nk}$ to match the shape of $\epsilon$ and $z_{nk}$. Additionally, we incorporate the cross-attention loss to strengthen the correlation between desired visual semantics and their corresponding pseudo-words, formulated as:
@@
when $A_{\theta}(p_{n},z_{nk}^{(t)})$ is the cross-attention maps between the pseudo-word $p_{n}$ and the noisy latent image $z_{nk}^{(t)}$ and $M_{nk}^{\prime\prime}$ is downsampled from $M_{nk}$ to match the shape of $A_{\theta}(p_{n},z_{nk}^{(t)})$.
Figure 4: Motivation of dynamic intensity. (a) Fixed intensity ($\alpha_{d}=0.5$ here) could cause noisy generated images. (b) Our dynamic intensity can mitigate noise memorization. (c) We report IQA results of Q-Align [47] on our dataset, showing that our dynamic intensity helps to enhance the quality of generated images.
Using $\mathcal{L}_{\text{diff}}$ and $\mathcal{L}_{\text{attn}}$, we first warm up the T2I model by jointly learning all samples, aiming to preliminarily inject the knowledge of visual semantics. The loss of the warm-up stage is defined as:
@@
where $\lambda_{\text{attn}}=0.01$ is the loss weight for $\mathcal{L}_{\text{attn}}$. For efficient fine-tuning, we only train the denoising U-Net $\epsilon_{\theta}$ in a low-rank adaptation (LoRA) [15] manner and the text embedding of the pseudo-words $\mathcal{P}$, keeping the others frozen. Thereafter, we employ Dual-Stream Balancing (DS-Bal) to address semantic imbalance. In this paradigm, the online denoising U-Net $\epsilon_{\theta}$ conducts sample-wise min-max optimization for the hardest-to-learn sample, and meanwhile the momentum denoising U-Net $\tilde{\epsilon}_{\theta}$ applies selective preserving regularization for the other samples.